OpenStudy (nosnip):

Write the equation of the line which passes through (8, 1) and is perpendicular to x = 4x-2

OpenStudy (nosnip):

and how can i work it out please?

OpenStudy (e.mccormick):

Do you know the relationship between the slope of a line and any line perpendicular to it?

OpenStudy (nosnip):

only that if it's a fraction the fraction is sort of flipped :/

OpenStudy (e.mccormick):

VERY CLOSE! You forgot the negative sign.

OpenStudy (nosnip):

okay, what do i do to work it out though?

OpenStudy (e.mccormick):

They are negative inverses. So if the slope of a line is \(m\) then the slope of a line perpendicular to it will be \(-\frac{1}{m}\) So if it is a fraction, it would be flipped and negative. That is where we start. What is the slope of your given line?

OpenStudy (nosnip):

4 so would the gradient be 1/4 or .25?

OpenStudy (nosnip):

-1/4 or -0.25 i mean

OpenStudy (e.mccormick):

YES! So, you have a slope and you were given the point (8,1). Do you know the point-slope formula?

OpenStudy (nosnip):

no

OpenStudy (e.mccormick):

Ah, that is the key here: \(y-y_1=m(x-x_1)\) Ever seen that?

OpenStudy (nosnip):

yes i think so

OpenStudy (e.mccormick):

Well, you take your point and set it to the sub one versions. So: \((8,1)=(x_1,y_1)\) Can you put those in the right place in the formula now?

OpenStudy (nosnip):

what would the other y and x be? i mean the ones that aren't x1 or y1?

OpenStudy (e.mccormick):

Just the letters.

OpenStudy (nosnip):

oh okay would it be y-1 = -0.25(x-8)

OpenStudy (e.mccormick):

Great! Now simplify it and get the y alone. Know what I mean?

OpenStudy (nosnip):

yep, give me a minute :)

OpenStudy (nosnip):

would it be y= -0.25-1

OpenStudy (e.mccormick):

What happned to the x? And you have \(-\frac{1}{4}\) distributing to -8... so not 1.

OpenStudy (e.mccormick):

Oh, or did you forget the sign on the -1 when you moved it from the left to the right?

OpenStudy (e.mccormick):

Hmmm... or some sign issue.

OpenStudy (nosnip):

don't you add 1 if i move it from one side to the other?

OpenStudy (e.mccormick):

Yah, but \(-\frac{1}{4}\times -8=??\)

OpenStudy (nosnip):

that would be -2 so i went from y-1 =-0.25x - 2 to y=-0.25x-1 because i added the 1

OpenStudy (e.mccormick):

Why would a negative times a negative be a negative?

OpenStudy (nosnip):

oh yes! i forgot the - on the 8 okay then it would be =3

OpenStudy (nosnip):

i mean +3

OpenStudy (e.mccormick):

/cheer

OpenStudy (nosnip):

so y = -0.25x+3

OpenStudy (e.mccormick):

Yes, or \(y=-\frac{1}{4}x+3\) if you like fractions. (Fractions are your friends! Really!)

OpenStudy (nosnip):

okay cheers :)

OpenStudy (e.mccormick):

The - sign is like pants. Embarrassing when you drop them at the wrong time. But once we got past that, it all came together.

OpenStudy (nosnip):

haha great thanks :)

OpenStudy (e.mccormick):

There is one other thing to be very careful of when you use the point slope formula. It is another sign problem. Want to see what I mean?

OpenStudy (nosnip):

yes please

OpenStudy (e.mccormick):

OK, what if I took the same line, but I wanted a perpendicualr at (-1,-6)?

OpenStudy (nosnip):

the minuses would make positives y+6 = -1/4(x+1) ?

OpenStudy (e.mccormick):

My \((x_1,y_1)=(-1,-6)\) So when I put them into \(y-y_1=m(x-x_1)\) I have to watch out for the sign! \(y-y_1=m(x-x_1)\implies (y-(-6)=m(x-(-1))\) and yah, you caught on.

OpenStudy (nosnip):

woohoo :)

OpenStudy (e.mccormick):

So that is the other sneaky sign problem to watch out for with these. If you can get down the formula, you should do fine.

OpenStudy (nosnip):

great thanks so much!

OpenStudy (e.mccormick):

no problem! Have fun!